1/1

Jean de La Fontaine,  Jean de La Fontaine,
The Cricket and the Ant  Die Grille und die Ameise
translated by Donald P. Webb  translated by Wilhelm Ernst Dohm

The cricket had sung her song all summer long but found her victuals too few when the north wind blew. Nowhere could she espy a single morsel of worm or fly. Her neighbor, the ant, might, she thought, help her in her plight, and she begged her for a little grain till summer would come back again. “By next August I’ll repay both Interest and principal; animal’s oath.” Now, the ant may have a fault or two But lending is not something she will do. She asked what the cricket did in summer. “By night and day, to any comer I sang whenever I had the chance.” “You sang, did you? That’s nice. Now dance.”

Grillchen, das den Sommer lang Zirpt’ und sang, Litt nun, da der Winter droht’, Harte Zeit und bittre Not: Nicht das kleinste Würmchen nur Und von Fliegen keine Spur! Und vor Hunger weinend leise, Schlich’s zur Nachbarin Ameise, Fleht’ sie an in ihrer Not, Ihr zu leihn ein Körnlein Brot, Bis der Sommer wiederkehre. “Glaub’ mir,” sprach’s, “auf Grillenehre, Vor dem Erntemond noch zahl’ Zins ich dir und Kapital.” Emschen, das, wie manche lieben Leute das Verleihen hasst, Fragt die Borgerin: “Was hast Du im Sommer denn getrieben?” - “Tag und Nacht hab’ ich ergötzt Durch mein Singen alle Leut’.” “Durch dein Singen? Sehr erfreut! Weißt du was? Dann tanze jetzt!”

Translation of the French poem “La Cigale et la Fourmi” by Jean de La Fontaine (∗1621-07.08 – †1695-04.12) into English by Donald P. Webb (∗1939).

Arg-549-1111 (2010-08-06 19:10:14)

Mi, Manfred Retzlaff, transprenis tiun ĉi tradukon el la retejo http: // www. bewilderingstories. com/ issue209/ cigale. html . La tradukinto, s-ro Don Webb, permesis tion mi. - I, Manfred Retzlaff, found this translation in the website http: // www. bewilderingstories. com/ issue209/ cigale. html . The author, Mr. Don Webb, gave me the permission.

Translation of the French poem “La Cigale et la Fourmi” by Jean de La Fontaine (∗1621-07.08 – †1695-04.12) into German by Wilhelm Ernst Dohm (Geburtsname: Elias Levy, ∗1819-05-24 – †1883-02-05).

Arg-549-1104 (2010-07-09 22:21:39)

Pri la tradukinto Wilhelm Ernst Dohm vidu la vikipedian retejon http: // de. wikipedia. org/ wiki/ Ernst_Dohm . Mi, Manfred Retzlaff, trovis tiun ĉi poem-germanigon en la libro “Jean de La Fontaine: Die Fabeln” (Gesamtausgabe in deutscher und französischer Sprache, Emil-Vollmer-Verlag, Wiesbaden.)

http://www.poezio.net/version?poem-id=549&version-id=1111,1104